### v1

# FroM: Frobenius Norm-Based Data-Free Adaptive Model Merging

Abstract
With the development of large language models, fine-tuning has emerged as an effective method to enhance performance in specific scenarios by injecting domain-specific knowledge. In this context, model merging techniques provide a solution for fusing knowledge from multiple fine-tuning models by combining their parameters. However, traditional methods often encounter task interference when merging full fine-tuning models, and this problem becomes even more evident in parameter-efficient fine-tuning scenarios. In this paper, we introduce an improvement to the RegMean method, which indirectly leverages the training data to approximate the outputs of the linear layers before and after merging. We propose an adaptive merging method called FroM, which directly measures the model parameters using the Frobenius norm, without any training data. By introducing an additional hyperparameter for control, FroM outperforms baseline methods across various fine-tuning scenarios, alleviating the task interference problem.
[…]
FroM: Frobenius Norm-Based Data-Free Adaptive Model Merging
Zijian Li  Xiaocheng Feng†‡  Huixin Liu Yichong HuangTing LiuBing Qin†‡ Harbin Institute of Technology  Peng Cheng Laboratory {lizijian, xcfeng, hxliu, ychuang, tliu, qinb}@ir.hit.edu.cn
1 Introduction
In recent years, significant advancements have been made in the fields of artificial intelligence and natural language processing, with large language models (LLMs) emerging as a focal point of attention Achiam et al. (2023); Guo et al. (2025). Among the various approaches, the pretraining and fine-tuning paradigm has emerged as the foundational framework for LLM technology Zhao et al. (2023); Han et al. (2021); Parthasarathy et al. (2024). With a growing number of fine-tuned downstream models, model merging has emerged as an effective approach to integrate models trained under different hyperparameter settings or for multi-task learning Jin et al. (2022). Moreover, model merging is commonly implemented in a data-free method, making it particularly advantageous for privacy-sensitive applications like federated learning McMahan et al. (2017); Wang et al. (2020).
Figure 1: Illustrative diagrams of fine-tuning and model merging.
Most existing model merging methods are based on linear operations such as weighted averaging Wortsman et al. (2022) and task arithmetic Ilharco et al. (2022). Weighted averaging combines parameters from multiple fine-tuned models, while task arithmetic enables multi-task learning and forgetting by adding and subtracting task vectors derived from parameter differences. Nevertheless, these methods encounter a fundamental challenge: simple linear combinations of parameters often lead to task interference Yadav et al. (2024). This issue is especially evident in parameter-efficient fine-tuning (PEFT) scenarios Ding et al. (2023), such as Low-Rank Adaptation (LoRA) Hu et al. (2021). Compared to full fine-tuning (FFT), the sparse structure of low-rank matrices in these methods further aggravates task interference Tang et al. (2023); Stoica et al. (2024). Figure 1 presents a schematic illustration of the fine-tuning and model merging processes.
To address this challenge, we take inspiration from the nonlinear weighted optimization method of RegMean Jin et al. (2022) and introduce a novel method, Frobenius Norm-Based Data-Free Adaptive Model Merging (FroM). Our method improves upon the RegMean method, which derives a closed-form solution by approximating the linear layer outputs of the models using the inner product matrix of the training data. However, RegMean still encounters challenges due to its reliance on access to training data, which may not always be available. To address this limitation, FroM adopts a data-free approach that directly measures model parameters using the Frobenius norm, and employs adaptive coefficients to adjust the discrepancies between task vectors. Our approach reduces the reliance on training data, while outperforming baselines in both full fine-tuning and LoRA scenarios, effectively mitigating the task interference issue.
Our main contributions are as follows:
  1. 1.
    We propose an adaptive model merging method, FroM, which utilizes the Frobenius norm to measure the differences between models to be merged, without any training data.
  2. 2.
    Compared to baseline methods, our approach is broadly applicable to FFT and PEFT models, such as LoRA, and achieves optimal or suboptimal results in various scenarios.
2 Related Work
Fine-tuning of LLMs.
Fine-tuning is a method for optimizing pre-trained LLMs for specific tasks. Through fine-tuning, LLMs not only excel in general tasks but can also meet the requirements of particular application scenarios Wei et al. (2021); Sanh et al. (2021); Chung et al. (2024). From the perspective of the amount of parameters updated, fine-tuning can be categorized into two approaches: FFT and PEFT. FFT updates all parameters of the pre-trained model, which often yields excellent performance on specific tasks. However, this approach is both time-consuming and computationally intensive. As models become larger and more complex, conducting FFT has become more challenging, particularly in environments with limited computational resources Lv et al. (2023). Alternatively, PEFT focuses on updating a small number of parameters to reduce computational costs and storage requirements, making it more widely used in resource-constrained settings or customized scenarios Xu et al. (2023b); Han et al. (2024). One of the most widely used methods within PEFT is LoRA Hu et al. (2021), which significantly reduces computational and storage costs by fine-tuning low-rank matrices, achieving highly effective fine-tuning with a limited number of parameters.
Model Merging.
Lu et al. (2024) introduce several collaboration strategies for LLMs, highlighting model merging as a technique that prevents models from getting trapped in local optima and improves multi-task collaboration ability. Traditional methods typically rely on weighted averaging of model parameters. Wortsman et al. (2022) introduce Model Soups, which apply both a weighted averaging method and a greedy merging strategy. Matena and Raffel (2022) adopt a different approach by utilizing the Fisher Information Matrix for parameter-level model merging, applying a weighted merging strategy to enhance model integration. Focusing on the outputs of linear layers, Jin et al. (2022) propose RegMean, a method that indirectly utilizes the training data through the inner product matrices to derive a closed-form solution for merging model parameters. Meanwhile, Ilharco et al. (2022) present a novel paradigm called task arithmetic, in which task vectors capture the ability shift of different fine-tuned models compared to the pre-trained model across various tasks. This allows for operations like forgetting and learning through arithmetic manipulation of the task vectors.
However, the task vectors of different models may inevitably lead to some degree of task interference. Prior research has demonstrated that weight disentanglement is crucial to the success of task arithmetic Ortiz-Jimenez et al. (2023). To address this, Yadav et al. (2024) introduce TIES-Merging, a strategy that minimizes redundant parameters and selects signs to reduce interference during model merging. In addition, the DARE method proposed by Yu et al. (2024) involves the random dropping of incremental parameters and rescaling operations. By integrating with other model merging algorithms, it enhances the performance of the model after fusion. According to the research by Tang et al. (2023); Stoica et al. (2024), existing merging methods are more challenging to apply to LoRA fine-tuned models compared to FFT models. Our method builds upon RegMean and introduces improvements that eliminate the need for obtaining the inner product matrix of training data. It can also be solved using a closed-form solution, supporting the integration of either FFT models or LoRA adapters.
3 Method
3.1 Motivation
The RegMean method naturally aims to make the output of the merged model close to the outputs of the original models. It cleverly uses the inner product matrix derived from the training data through a closed-form solution, while avoiding data privacy issues that may arise from directly using the training data. However, this brings about additional requirements for model release, as open-source models typically do not provide the inner product matrix. This limitation significantly impacts the applicability of the RegMean method.
Therefore, we propose a different approach: instead of measuring the output discrepancies between the merged model and the original models, why not directly measure the model weights themselves? However, directly measuring model weights is equivalent to the weighted averaging method, which has limited effectiveness. We draw inspiration from research on model pruning, where compression is achieved by removing connections with small weights Han et al. (2015). This aligns with our intuition that larger task vectors generally contain more information. We adopt a holistic, layer-wise metric based on the squared Frobenius norm, which measures the overall magnitude of weight matrices and serves as an effective indicator of model differences. For this reason, we use the Frobenius norm of the task vector to quantify model differences and raise it to the power of $k$ to serve as a weighting coefficient, with $k$ as a hyperparameter controlling this balance.
3.2 Merging of FFT Models
We use the objective function in Equation 1 to measure the performance difference between the task vectors before and after merging.
$\mathcal{L}(\boldsymbol{\theta})=\sum_{i=1}^{n}\|\boldsymbol{\theta}_{i}\|_{F}^{k}\|\boldsymbol{\theta}-\boldsymbol{\theta}_{i}\|_{F}^{2}$ (1)
where $n$ denotes the number of fine-tuned models, $\boldsymbol{\theta}_{i}$ represents the task vector of the $i$-th model, $\boldsymbol{\theta}$ is the task vector after merging, and $k$ is a hyperparameter. The comparison between the RegMean and FroM Methods is illustrated in Figure 2.
Since the above function is convex, a closed-form solution (i.e., the point where the gradient is zero) can be derived. The objective function can be restructured, resulting in:
$\mathcal{L}(\boldsymbol{\theta})$ $=\sum_{i=1}^{n}\|\boldsymbol{\theta}_{i}\|_{F}^{k}\operatorname{Tr}\left((\boldsymbol{\theta}-\boldsymbol{\theta}_{i})^{\top}(\boldsymbol{\theta}-\boldsymbol{\theta}_{i})\right)$ (2)
$=\sum_{i=1}^{n}\|\boldsymbol{\theta}_{i}\|_{F}^{k}\operatorname{Tr}\left(\boldsymbol{\theta}^{\top}\boldsymbol{\theta}-2\boldsymbol{\theta}^{\top}\boldsymbol{\theta}_{i}+\boldsymbol{\theta}_{i}^{\top}\boldsymbol{\theta}_{i}\right)$ (3)
The last term in the parentheses is a constant; therefore, the derivative with respect to $\boldsymbol{\theta}$ is:
$\frac{\partial\mathcal{L}(\boldsymbol{\theta})}{\partial\boldsymbol{\theta}}=2\sum_{i=1}^{n}\|\boldsymbol{\theta}_{i}\|_{F}^{k}\boldsymbol{\theta}-2\sum_{i=1}^{n}\|\boldsymbol{\theta}_{i}\|_{F}^{k}\boldsymbol{\theta}_{i}$ (4)
By setting $\frac{\partial\mathcal{L}(\boldsymbol{\theta})}{\partial\boldsymbol{\theta}}=0$, the closed-form solution is derived, as shown in Equation 5. In fact, we observe that as $k\to 0$, the final result corresponds to the simple averaging method, whereas as $k\to+\infty$, the result converges to the method that selects the weight with the largest norm. Therefore, this approach can be viewed as a balance between these two methods.
$\boldsymbol{\theta}^{*}=\frac{\sum_{i=1}^{n}\|\boldsymbol{\theta}_{i}\|_{F}^{k}\boldsymbol{\theta}_{i}}{\sum_{i=1}^{n}\|\boldsymbol{\theta}_{i}\|_{F}^{k}}$ (5)
Figure 2: Comparison between the RegMean and FroM Methods, where $X$ represents the input to the linear layer.
3.3 Merging of LoRA
The task arithmetic-based approach poses challenges in integrating LoRA Tang et al. (2023); Stoica et al. (2024), making it difficult both to merge and to control the rank of the merged model. To address this issue, we modify the objective function based on Equation 1, incorporating LoRA-type structures while measuring weight differences using the Frobenius norm. The revised objective function is given in Equation 6, where $\theta_{i}$ refers to the task vector in its original form before being flattened into a one-dimensional vector. When merging parameters layer-wise across different models, we employ LoRA-type structures, specifically utilizing the $\mathbf{A}$ and $\mathbf{B}$ matrices as components optimized in a manner similar to the LoRA mechanism. Moreover, by explicitly specifying the rank of $\mathbf{A}$ and $\mathbf{B}$ (i.e., the row count of $\mathbf{A}$ and the column count of $\mathbf{B}$), we enable dynamic rank adjustment. This approach not only increases flexibility in model merging but also facilitates a certain degree of compression during model storage.
$\mathcal{L}(\mathbf{A},\mathbf{B})=\sum_{i=1}^{n}\|\boldsymbol{\theta}_{i}\|_{F}^{k}\|\mathbf{B}\mathbf{A}-\boldsymbol{\theta}_{i}\|_{F}^{2}$ (6)
Since Equation 6 is a non-convex function, it may have multiple local minima. Therefore, an alternating optimization algorithm is required to solve it. Specifically, by initializing $\mathbf{A}$ from a normal distribution with mean 00 and variance $\sigma^{2}$, and treating either $\mathbf{A}$ or $\mathbf{B}$ as an invariant, we iteratively update $\mathbf{B}$ and $\mathbf{A}$ using Equation 7 and Equation 8, respectively. In these equations, $\dagger$ denotes the Moore-Penrose inverse (see Appendix A for the proof). This procedure for approximating a local minimum of the objective function is shown in Algorithm 1. Due to the potential computational errors in the calculation of the MP inverse, an early stopping mechanism is implemented, which halts the iteration when the loss begins to increase.
$\mathbf{B}=\left(\sum_{i=1}^{n}\|\boldsymbol{\theta}_{i}\|_{F}^{k}\boldsymbol{\theta}_{i}\mathbf{A}^{\top}\right)\left(\sum_{i=1}^{n}\|\boldsymbol{\theta}_{i}\|_{F}^{k}\mathbf{A}\mathbf{A}^{\top}\right)^{\dagger}$ (7)
$\mathbf{A}=\left(\sum_{i=1}^{n}\|\boldsymbol{\theta}_{i}\|_{F}^{k}\mathbf{B}^{\top}\mathbf{B}\right)^{\dagger}\left(\sum_{i=1}^{n}\|\boldsymbol{\theta}_{i}\|_{F}^{k}\mathbf{B}^{\top}\boldsymbol{\theta}_{i}\right)$ (8)
3.4 Complexity
We next analyze the computational complexity of the two merging methods discussed above. We let $n$ denote the number of models to be merged, and $d$ represent the number of parameters in the FFT models. Therefore, the complexity of merging FFT models is $O(nd)$. For the merging of LoRA (i.e., Algorithm 1), we simplify the analysis by assuming that LoRA is applied to every layer (in practice, it is typically applied to only a subset of layers). We let $L$ denote the number of layers in the model and $T$ the number of iterations, assuming that each layer’s weight matrix has dimensions $(d_{1},d_{2})$, with the corresponding LoRA matrices $B$ and $A$ are of sizes $(d_{1},r)$ and $(r,d_{2})$, respectively. The first term in Equation 7 involves matrix multiplication with complexity $O(d_{2}^{2}r)$. The second term involves both matrix multiplication and matrix inversion, with respective complexities $O(d_{2}r^{2})$ and $O(r^{3})$. Thus, the overall complexity becomes: $O(d_{2}^{2}r+d_{2}r^{2}+r^{3})=O(d_{2}^{2}r)$, since $r\ll min(d_{1},d_{2})$. Similarly, the complexity of Equation 8 is $O(d_{1}^{2}r)$. Therefore, the overall complexity of the algorithm is: $O(nLTr(d_{1}^{2}+d_{2}^{2}))$.
4 Experimental Setup
Algorithm 1 Alternating Optimization-Based Model Merging Algorithm Utilizing LoRA-Type Structures
1:  Initialize matrix $\mathbf{A}\sim\mathcal{N}(0,\sigma^{2})$
2:  Initialize early stopping flag early_stopping
3:  Set iteration count iters
[…]
5:  for $i=0$ to $\texttt{iters}-1$ do
6:     Compute matrix $\mathbf{B}$ using Equation 7
7:     Compute matrix $\mathbf{A}$ using Equation 8
8:     if early_stopping then
9:        Compute loss using Equation 6
10:        if loss < min_loss then
[…]
12:        else
13:           break
14:        end if
15:     end if
16:  end for
17:  Save matrices $\mathbf{A}$ and $\mathbf{B}$
Merging of FFT Models.
We test two different model architectures, each based on the same foundational model, across three distinct FFT tasks: Instruction-Following, Mathematical Reasoning, and Code-Generating. First, we adopt the experimental setup of Yu et al. (2024), utilizing models based on Llama-2-13B Touvron et al. (2023): WizardLM-13B Xu et al. (2023a), WizardMath-13B Luo et al. (2023), and llama-2-13b-code-alpaca [^1] https://huggingface.co/layoric/llama-2-13b-code-alpaca. Second, we utilize models based on Qwen2.5-7B [^2] https://huggingface.co/Qwen/Qwen2.5-7B Yang et al. (2024): Qwen2.5-7B-Instruct [^3] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct, Qwen2.5-Math-7B-Instruct [^4] https://huggingface.co/Qwen/Qwen2.5-Math-7B-Instruct, and Qwen2.5-Coder-7B-Instruct [^5] https://huggingface.co/Qwen/Qwen2.5-Coder-7B-Instruct. For testing, we employ the Task Arithmetic Ilharco et al. (2022), TIES-Merging Yadav et al. (2024), and our FroM methods, additionally incorporating the DARE method Yu et al. (2024) in combination with TIES-Merging and FroM for comparative evaluation. We do not include the RegMean method in our comparison, as we are unable to obtain the inner product matrices of all model training datasets, which highlights a significant practical limitation of RegMean. For both evaluation settings involving different base models, we set the linear weighting coefficient $\alpha$ to $1.0$ (see Appendix B.2 for details), and use $k=1.0$ in our FroM method. For the choice of the hyperparameter $k$ in practical applications, refer to Section 5.3.
Regarding benchmark selection, we follow the setup of Yu et al. (2024), utilizing AlpacaEval Li et al. (2023) to assess the models’ instruction-following abilities, GSM8K Cobbe et al. (2021) and MATH Hendrycks et al. (2021) to evaluate their mathematical reasoning skills, and HumanEval Chen et al. (2021) and MBPP Austin et al. (2021) for testing code generation capabilities. Further details can be found in Appendix B.
Merging of LoRA.
We fine-tune a binary classification head on top of the Meta-Llama-3-8B [^6] https://huggingface.co/meta-llama/Meta-Llama-3-8B model. First, we initialize the classifier by training it on the MNLI dataset from the GLUE benchmark Wang et al. (2019). Then, we fine-tune the model using the binary text entailment datasets from the GLUE benchmark, specifically QNLI, RTE, and WNLI, with the LoRA approach. For each task, we train the model for 10 epochs and select the best checkpoint for evaluation. We not only test the merging of LoRA across three tasks but also examine the merging of the three optimal checkpoints for the same task. We apply the Task Arithmetic, RegMean, TIES-Merging, DARE method (which combines the TIES-Merging approach), KnOTS method (which integrates the first three methods) Stoica et al. (2024), and FroM method (with $k=0.9$). The linear weighting coefficient is set to $\alpha=0.7$, determined via a linear search. Further details, including the search results and relevant settings, are provided in Appendix B.
5 Main Results
5.1 Merging of FFT Models
Models Merging Methods
Instruction-
Following
Mathematical
Reasoning
Code-Generating Avg.
AlpacaEval GSM8K MATH HumanEval MBPP
LM / 51.09 45.11 5.92 32.93 31.20 33.25
Math / / 59.97 11.60 / / 35.78
Code / / / / 24.39 28.00 26.20
LM & Math Task Arithmetic 50.55 41.62 6.56 / / 32.91
TIES-Merging 50.42 51.55 7.22 / / 36.40
[…]
LM & Code Task Arithmetic 53.23 / / 32.93 30.60 38.92
TIES-Merging 46.33 / / 0.00 0.00 15.44
DARE+TIES-Merging 47.67 / / 0.00 0.00 15.89
[…]
Math & Code Task Arithmetic / 58.45 12.32 5.49 7.00 20.82
TIES-Merging / 58.00 12.28 9.15 19.00 24.61
[…]
LM & Math & Code Task Arithmetic 52.84 42.00 6.20 15.85 18.20 27.02
TIES-Merging 46.24 47.08 5.32 0.00 0.00 19.73
[…]
DARE+FroM 49.12 55.34 8.06 29.09 31.80 34.68
Table 1: Merging results of WizardLM-13B, WizardMath-13B, and llama-2-13b-code-alpaca models based on Llama-2-13B. Bold and underlined text indicate the optimal and suboptimal results, respectively.
Models Merging Methods
Instruction-
Following
Mathematical
Reasoning
Code-Generating Avg.
AlpacaEval GSM8K MATH HumanEval MBPP
LM / 16.24 79.08 14.40 79.88 65.40 51.00
Math / / 75.74 1.12 / / 38.43
Code / / / / 30.49 39.00 34.74
LM & Math Task Arithmetic 2.25 18.95 6.04 / / 9.08
TIES-Merging 0.00 0.68 0.00 / / 0.23
DARE+TIES-Merging 0.00 0.15 0.00 / / 0.05
[…]
DARE+FroM 0.00 0.45 0.00 / / 0.15
[…]
TIES-Merging 0.00 / / 0.00 0.00 0.00
DARE+TIES-Merging 0.00 / / 0.00 0.00 0.00
[…]
DARE+FroM 0.00 / / 0.00 0.00 0.00
Math & Code Task Arithmetic / 0.38 0.00 0.00 0.00 0.09
[…]
DARE+TIES-Merging / 0.38 0.00 0.00 0.00 0.09
FroM / 17.36 0.00 1.22 0.80 4.85
DARE+FroM / 0.38 0.00 0.00 0.00 0.09
LM & Math & Code Task Arithmetic 0.00 0.45 0.00 0.00 0.00 0.09
TIES-Merging 0.00 0.08 0.00 0.00 0.00 0.02
DARE+TIES-Merging 0.00 0.23 0.00 0.00 0.00 0.05
[…]
Table 2: Merging results of Qwen2.5-7B-Instruct, Qwen2.5-Math-7B-Instruct, and Qwen2.5-Coder-7B-Instruct models based on Qwen2.5-7B. Bold and underlined text indicate the optimal and suboptimal results, respectively.
The merging results of the WizardLM-13B, WizardMath-13B, and llama-2-13b-code-alpaca models based on Llama-2-13b are shown in Table 1. As observed, for the merging of the first two models, our FroM method achieves better average performance, outperforming the previous optimal baseline by 0.73%. Similarly, when merging WizardLM-13B and llama-2-13b-code-alpaca, FroM also achieves the highest average score. For the merging of the latter two models, however, the Task Arithmetic and TIES-Merging methods demonstrate better performance. In the case of merging three models, the Task Arithmetic method performs best on the AlpacaEval benchmark, but shows moderate performance across the other four datasets. For tasks related to Mathematical Reasoning and Code-Generating, both our FroM method and its integration with DARE exhibit strong performance, with the latter achieving the best average performance, surpassing the previous optimal baseline by 1.68%. Therefore, although the FroM method does not consistently outperform the baseline across every benchmark, a careful selection of the hyperparameter $k$ enables a well-balanced trade-off, ultimately leading to superior performance compared to other baseline methods.
For the base model Qwen2.5-7B, we present the results of merging the models Qwen2.5-7B-Instruct, Qwen2.5-Math-7B-Instruct, and Qwen2.5-Coder-7B-Instruct, as shown in Table 2. Unexpectedly, the performance of the merged models deteriorates significantly, except when merging Qwen2.5-7B-Instruct and Qwen2.5-Coder-7B-Instruct using the Task Arithmetic or FroM methods, which yield comparatively better results. When the other two models are integrated with Qwen2.5-Math-7B-Instruct, a substantial decline in performance is observed. Based on the specific test outputs, we observe that the majority of the model’s responses are either incoherent or irrelevant to the given task. This suggests a significant degree of task interference between Qwen2.5-Math-7B-Instruct and the other two models. After extensive fine-tuning on different datasets, the linear mode connectivity Ainsworth et al. (2022); Zhou et al. (2023) among these models tends to be disrupted, leading to the failure of linear weighting methods. This finding indicates that not all fine-tuned models derived from the same base model are equally suitable for model merging. Furthermore, the merging results using the TIES-Merging and DARE methods are close to zero (with outputs consisting entirely of garbled text), indicating that these methods aggravate task interference. In contrast, our FroM method demonstrates greater robustness when merging both two and three models.
5.2 Merging of LoRA
For the LoRA merging experiments, we train a shared classification head on the MNLI dataset. The performance of this classification head on other datasets is presented in the first row of Table 3. The accuracy across different tasks exceeds 50%, demonstrating a reasonable degree of generalization, considering that for binary classification tasks, the probability of randomly selecting the correct answer is 50%. We further fine-tune on the QNLI, RTE, and WNLI datasets, selecting the best-performing checkpoints for subsequent experiments.
The testing results for different LoRA merging methods are summarized in Table 3. It is evident that the application of the TIES-Merging and DARE methods results in a decline in model accuracy, indicating that these approaches may not be universally effective across all downstream tasks and may perform particularly poorly on simpler ones. On the other hand, our FroM method continues to demonstrate effectiveness, with its average performance surpassing other baselines. Notably, for the QNLI task, our method even exceeds the baseline accuracy, highlighting that our approach enhances the generalization capability of the merged model resulting from multi-task learning.
Merging Methods QNLI RTE WNLI Avg.
Baseline
/ 53.74 59.21 54.93 55.96
Fine-tuning
QNLI 94.05 56.32 43.66 64.68
RTE 54.75 71.49 64.79 63.68
WNLI 54.26 62.82 54.69 57.26
3-Model Merging
Task Arithmetic 90.10 67.51 52.11 69.91
RegMean 69.84 66.34 53.52 63.23
TIES-Merging 65.50 65.70 61.97 64.39
DARE+TIES-Merging 54.26 60.65 56.34 57.08
KnOTS+Task Arithmetic 94.44 66.43 50.70 70.52
KnOTS+TIES-Merging 93.65 64.98 52.11 70.25
KnOTS+DARE+TIES-Merging 59.07 63.54 60.56 61.06
[…]
Table 3: Test results of three models fine-tuned using LoRA on QNLI, RTE, and WNLI datasets with different merging methods. Bold text indicates the optimal results.
We also test the merging of the three best checkpoints from the same dataset. As shown in Figure 4, our FroM method achieves the best performance, surpassing the accuracy of the original fine-tuning across all three datasets. For comparison, the original KnOTS method performs comparably to our approach, while Task Arithmetic yields significantly inferior results. The other methods demonstrate relatively moderate performance. This further demonstrates the effectiveness of our method in merging checkpoints for the same task.
Figure 3: Results of merging WizardLM-13B, WizardMath-13B, and LLaMA-2-13B-Code-Alpaca models with different values of $k$. The optimal results before merging are represented by the gray dashed line.
5.3 Ablation Study
For the WizardLM-13B, WizardMath-13B, and llama-2-13b-code-alpaca models, we test the effect of varying the value of $k$ in Equation 1. Specifically, we select $k=0$ and integer powers of 2 ranging from $2^{-3}$ to $2^{3}$, and plot the results under these different conditions. As shown in Figure 3, the performance is generally poor when $k\to 0$ or $k\to+\infty$. The best results occur when $k$ is set to 0.5 or 1.0, with accuracy on all benchmarks closely matching the optimal performance of the fine-tuned models.
We also test the effect of varying the value of $k$ on LoRA fusion, as shown in Figure 5. The optimal average accuracy is achieved when $k$ is set to 1.0. As the value of $k$ increases, the performance of the fused model improves on the QNLI dataset but decreases on the RTE and WNLI datasets. This change can be attributed to the larger size of the QNLI dataset, which results in greater influence of its fine-tuned model parameters during merging.
Figure 4: Accuracy comparison after merging the three optimal LoRA checkpoints on QNLI, RTE, and WNLI datasets.
6 Conclusion
In this paper, we propose a novel merging approach called FroM, which utilizes the Frobenius norm to quantify weight discrepancies between models without any training data. This method adjusts adaptively through a hyperparameter, and the merged parameters are derived using a closed-form solution. For the experimental part, we first test our method on three models based on Llama-2-13B, and the results show that FroM outperforms existing baseline methods. Next, while baseline methods struggle to effectively integrate three fine-tuned models based on Qwen2.5-7B, our approach outperforms the baseline methods, demonstrating its effectiveness and robustness. Finally, when merging LoRA adapters, FroM also shows superior performance over other methods. Overall, the proposed FroM method demonstrates outstanding performance in both FFT and LoRA scenarios, effectively alleviating the task interference issue in model merging and showcasing strong applicability. We hope this work inspires future research in the model merging field, helps address existing task interference issues, and provides a new perspective for tackling these challenges.
Figure 5: Results of merging the three LoRAs. The average accuracy is depicted by a dashed line.
Limitations
The limitations of our FroM method can be summarized as follows: (1) the necessity for a more comprehensive theoretical analysis of the algorithm, and (2) the crucial selection of the hyperparameter $k$. Future research directions with significant potential include providing an interpretable metric for task interference. By analyzing the upper bound of the model’s performance, it may be possible to assess and determine whether the model merging process is feasible.
Appendix
Appendix A Proof of the expression
The objective function given in Equation 6 can be reformulated as follows:
$\mathcal{L}(\mathbf{A},\mathbf{B})$ $=\sum_{i=1}^{n}\|\boldsymbol{\theta}_{i}\|_{F}^{k}\|\mathbf{B}\mathbf{A}-\boldsymbol{\theta}_{i}\|_{F}^{2}$
$=\sum_{i=1}^{n}\|\boldsymbol{\theta}_{i}\|_{F}^{k}\operatorname{Tr}\left((\mathbf{B}\mathbf{A}-\boldsymbol{\theta}_{i})^{\top}(\mathbf{B}\mathbf{A}-\boldsymbol{\theta}_{i})\right)$
$=\sum_{i=1}^{n}\|\boldsymbol{\theta}_{i}\|_{F}^{k}\operatorname{Tr}\left(\mathbf{A}^{\top}\mathbf{B}^{\top}\mathbf{B}\mathbf{A}-\boldsymbol{\theta}_{i}^{\top}\mathbf{B}\mathbf{A}-\mathbf{A}^{\top}\mathbf{B}^{\top}\boldsymbol{\theta}_{i}+\boldsymbol{\theta}_{i}^{\top}\boldsymbol{\theta}_{i}\right)$
$=\sum_{i=1}^{n}\|\boldsymbol{\theta}_{i}\|_{F}^{k}\operatorname{Tr}(\mathbf{A}^{\top}\mathbf{B}^{\top}\mathbf{B}\mathbf{A})-2\sum_{i=1}^{n}\|\boldsymbol{\theta}_{i}\|_{F}^{k}\operatorname{Tr}(\mathbf{A}^{\top}\mathbf{B}^{\top}\boldsymbol{\theta}_{i})+\sum_{i=1}^{n}\|\boldsymbol{\theta}_{i}\|_{F}^{k}\operatorname{Tr}\left(\boldsymbol{\theta}_{i}^{\top}\boldsymbol{\theta}_{i}\right)$
Treating matrix $\mathbf{A}$ as a constant, we differentiate matrix $\mathbf{B}$ with respect to it, noting that the final term is a constant, which yields:
$\frac{\partial\mathcal{L}(\mathbf{A},\mathbf{B})}{\partial\mathbf{B}}=2\sum_{i=1}^{n}\|\boldsymbol{\theta}_{i}\|_{F}^{k}\mathbf{B}\mathbf{A}\mathbf{A}^{\top}-2\sum_{i=1}^{n}\|\boldsymbol{\theta}_{i}\|_{F}^{k}\boldsymbol{\theta}_{i}\mathbf{A}^{\top}$
Let $\frac{\partial\mathcal{L}(\mathbf{A},\mathbf{B})}{\partial\mathbf{B}}=0$, then:
$\mathbf{B}=\left(\sum_{i=1}^{n}\|\boldsymbol{\theta}_{i}\|_{F}^{k}\boldsymbol{\theta}_{i}\mathbf{A}^{\top}\right)\left(\sum_{i=1}^{n}\|\boldsymbol{\theta}_{i}\|_{F}^{k}\mathbf{A}\mathbf{A}^{\top}\right)^{\dagger}$
Similarly, treating matrix $\mathbf{B}$ as a constant, we differentiate with respect to matrix $\mathbf{A}$, yielding:
$\frac{\partial\mathcal{L}(\mathbf{A},\mathbf{B})}{\partial\mathbf{A}}=2\sum_{i=1}^{n}\|\boldsymbol{\theta}_{i}\|_{F}^{k}\mathbf{B}^{\top}\mathbf{B}\mathbf{A}-2\sum_{i=1}^{n}\|\boldsymbol{\theta}_{i}\|_{F}^{k}\mathbf{B}^{\top}\boldsymbol{\theta}_{i}$
Let $\frac{\partial\mathcal{L}(\mathbf{A},\mathbf{B})}{\partial\mathbf{A}}=0$, then:
$\mathbf{A}=\left(\sum_{i=1}^{n}\|\boldsymbol{\theta}_{i}\|_{F}^{k}\mathbf{B}^{\top}\mathbf{B}\right)^{\dagger}\left(\sum_{i=1}^{n}\|\boldsymbol{\theta}_{i}\|_{F}^{k}\mathbf{B}^{\top}\boldsymbol{\theta}_{i}\right)$
Appendix B Experimental Details
B.1 Fine-tuning Settings
For the integration of LoRA adapters, the hyperparameters are set as follows: the ranks of matrices A and B are fixed at 16, the batch size is set to 32, and the learning rate is set to $51e-$ 5. We first train a classification head using the MNLI dataset. Since MNLI is a three-class task, we modify it to a binary classification output by combining the neutral and contradiction categories into a single not_entailment class. For each task within the QNLI, RTE, and WNLI datasets, we train the models for 10 epochs and select the checkpoint with the highest accuracy for subsequent LoRA adapter integration.
B.2 Test Details
For the integration of FFT models, we perform a linear search over the linear weighting coefficient $\alpha$ used to merge the three Llama-2-13B-based models. The results are presented in Table 4. Based on these results, we select $\alpha=1.0$ as the final linear weighting coefficient to achieve a more balanced outcome. For the Qwen2.5-7B-based models, due to their poor merging performance, we also adopt $\alpha=1.0$ without extensive tuning.
Regarding the evaluation, we conducted tests using AlpacaEval, GSM8K, MATH, HumanEval, and MBPP datasets. In the case of AlpacaEval, the model’s performance is evaluated using the gpt-3.5-turbo model. Additionally, the final success rate metric selected is length_controlled_winrate, rather than win_rate. When merging models from the Qwen-2.5 series, we minimized the use of line breaks in the prompts. Excessive line breaks negatively impact the model’s responses, leading to the generation of repetitive and meaningless text.
When merging LoRA adapters, we also employ a linear search strategy to determine the optimal linear weighting coefficient $\alpha$ and the hyperparameter $k$ used in our FroM method. Based on this search, we select $\alpha=0.7$ and $k=0.9$. Detailed results are presented in Table 5 and  6.
$\alpha$ AlpacaEval GSM8K MATH HumanEval MBPP Avg.
0.5 49.76 61.79 10.32 6.71 4.60 26.64
0.6 53.88 63.53 9.60 8.54 3.00 27.71
0.7 55.35 62.40 7.92 10.37 7.20 28.65
0.8 56.58 60.05 7.18 11.59 13.80 29.84
0.9 57.00 55.57 5.62 4.88 17.00 28.01
1.0 52.84 42.00 6.20 15.85 18.20 27.02
Table 4: Results of the linear search for the linear weighting coefficient used in merging Llama-2-13B-based models.
$\alpha$ QNLI RTE WNLI Avg.
0.5 70.18 68.23 53.52 63.98
0.6 81.59 67.87 57.75 69.07
0.7 90.10 67.51 52.11 69.91
0.8 93.78 66.06 49.30 69.71
0.9 94.84 64.98 46.48 68.77
1.0 78.69 68.23 57.75 68.22
Table 5: Results of the linear search for the linear weighting coefficient used in merging LoRA adapters.
$k$ QNLI RTE WNLI Avg.
0.5 92.00 65.34 50.70 69.35
0.6 92.84 66.06 49.30 69.40
0.7 93.54 64.98 52.11 70.21
0.8 94.07 64.62 53.52 70.74
0.9 94.51 64.26 53.52 70.76
1.0 94.82 63.18 52.11 70.04
Table 6: Results of selecting different values of $k$ in LoRA merging for our FroM method.